# Memorization to Generalization: Emergence of Diffusion Models from Associative Memory

### v2

@@ Abstract
Hopfield networks are associative memory (AM) systems, designed for storing and retrieving patterns as local minima of an energy landscape. In the classical Hopfield model, an interesting phenomenon occurs when the amount of training data reaches its critical memory load — spurious states, or unintended stable points, emerge at the end of the retrieval dynamics, leading to incorrect recall. In this work, we examine diffusion models, commonly used in generative modeling, from the perspective of AMs. The training phase of diffusion model is conceptualized as memory encoding (training data is stored in the memory). The generation phase is viewed as an attempt of memory retrieval. In the small data regime the diffusion model exhibits a strong memorization phase, where the network creates distinct basins of attraction around each sample in the training set, akin to the Hopfield model below the critical memory load. In the large data regime, a different phase appears where an increase in the size of the training set fosters the creation of new attractor states that correspond to manifolds of the generated samples. Spurious states appear at the boundary of this transition and correspond to emergent attractor states, which are absent in the training set, but, at the same time, have distinct basins of attraction around them. Our findings provide: a novel perspective on the memorization-generalization phenomenon in diffusion models via the lens of AMs, theoretical prediction of existence of spurious states, empirical validation of this prediction in commonly-used diffusion models.The code is available at: https://github.com/Lemon-cmd/Diffusion-Models-and-Associative-Memory denotes equal contribution.
@@ 1 Introduction
Hopfield networks are the simplest energy-based models that conceptualize memories as attractor states corresponding to local minima of their energy function, and the memory retrieval as dynamical convergence towards those attractor states [1, 2, 3]. Recently, they have seen a resurgence of interest due to advances in their memorization capacity. Notably, Dense Associative Memories (DenseAMs), which are extensions of Hopfield networks with super-linear memory capacity [4, 5] have paved the way for more sophisticated associative memory systems [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], driven in part by their connection to the attention mechanism in transformers [19, 20, 13] and to neurobiology [20, 21]. Simultaneously, generative diffusion models [22] have gained considerable popularity, due to their flexibility and accuracy in modeling high-dimensional distributions for a variety of domains —such as image generation [23, 24, 25], audio [26, 27, 28], video synthesis [29, 30, 31, 32], and other scientific applications. However, despite their effectiveness, diffusion models pose challenges related to privacy and security, as concerns grow about their tendency to replicate training data [33, 34, 35]. Such matters consequently emphasize the need for further understanding of memorization and generalization behaviors in diffusion models.
Recent works [36, 37, 38] have begun establishing theoretical connections between DenseAMs and generative diffusion models, offering a foundation for bridging the two fields. Specifically, it was shown that the logarithm of the probability of the generated samples in diffusion models can be interpreted as the energy function of a commonly used DenseAM model with the softmax activation function. This makes it possible to apply theoretical tools developed for associative memory to better understand the computational properties of diffusion models.
A possible key that can further bridge these two systems and highlight the properties of the memorization-to-generalization transition in diffusion models is the phenomenon of spurious patterns. Historically considered as detrimental to pattern recall [39, 40, 41], spurious patterns can be interpreted as combinations or interpolations of stored patterns, hinting at the network’s ability to synthesize new patterns from existing training data. This blending of fundamental memories resembles the generalization process in generative models, where learned representations are used to generate novel outputs. In this way, spurious patterns offer a fascinating framework for exploring the balance between memorization — where models store exact patterns from training data — and generalization, where they use underlying features to create genuinely new samples [42]. Studying the conditions under which spurious patterns emerge and how they contribute to the network’s behavior can thus shed light on generalization in both associative memory and contemporary generative models.
Previous works have explored memorization in generative models through various approaches. For instance, [43, 44] proposed general methods to measure memorization, while [33, 34, 45, 46] examined the memorization capacity in diffusion models as a function of the size of the training data. Other works have focused on understanding generalization, such as the theoretical estimates of the generalization gap [47], and the spectral analyses [48, 49, 50]. A recent study [51] proposed that generalization in diffusion models arises as a result of the combination of limited capacity and structural constraints neural network, which results in predictable patterns at the boundary between memorization and generalization. Other recent studies [52, 50, 49] explored memorization and generalization through the relationship between the dimensionalities of the ground-truth manifold and the manifold that is learned by the model. Meanwhile, [53] explored the evolution of the generated density as a function of the training time to provide an analysis on how mode collapse occurs and leads to the loss of diversity in the generated density. Finally, [54] and [55] characterized memorization in diffusion models as “condensation” phase transitions known in the statistical physics of spin glasses and random energy models. Although these studies shed some light on memorization and generalization, they do not fully explore the model’s behavior during the transition between these two phases.
Figure 1: A simple illustration depicting the change in the energy landscape as the size of the training dataset is increased. In the small data regime, the model memorizes the training data points as local minima of the energy. When the amount of training data exceeds the memory capacity of the model, spurious patterns are formed and training data points are no longer energy minima. Subsequent increase of the training set size leads to the generalization phase, which is defined by the formation of continuous manifold of the low energy states.
@@
Contributions.   Most of the mentioned works on memorization-to-generalization transition tackled it in a generalization-centric fashion. Namely, the memorization phenomenon is typically viewed as a “small side effect” of the diffusion models. In our work we adopt a complimentary approach. From the beginning, we cast the diffusion modeling pipeline into the associative memory framework. The training phase of the diffusion model is conceptualized as writing into the memory operation. The generation phase is viewed as an attempt of memory recall. This recall can be successful, resulting in the retrieval of training samples (memorization), or unsuccessful, resulting in the generation of new previously unseen samples (generalization). Thus, according to this theory, generalization (creation of genuinely novel samples) is a failure of memory recall. This point of view allows us to apply theory developed in DenseAMs for analyzing the memorization-to-generalization transition.
In DenseAMs, it is well-known that successful memory recall deteriorates as the number of the stored data points increases [4]. Importantly, the transition from successful to unsuccessful memory recall is marked by the peculiar phenomenon — the emergence of spurious states, see Figure 1. Our theory provides a specific and empirically verifiable prediction — spurious states must exist in the conventional diffusion models, which are trained and ran using standard methods. In this paper we demonstrate that this is indeed the case. Spurious states emerge at the boundary between the memorization and generalization phases, see examples of spurious samples in Figure 2.
@@
In accord with previous literature, we find that (1) diffusion models undergo a transition from memorization to generalization as the number of training points increases; (2) unlike previous works on diffusion models, at the onset of generalization we detect the emergence of spurious patterns, which have sizable basins of attraction, but are not memorized data points; and lastly, (3) we provide theoretical descriptions distinguishing these spurious states from memorized and generalized patterns in terms of energy landscapes. All of these findings suggest a close parallel between diffusion models and DenseAMs above the critical storage capacity.
Figure 2: Examples of memorized, spurious, and generalized samples in their respective columns for four datasets. For each target image (shown on the left), its top-4 nearest neighbors from the training set (top row) and the synthetic set (bottom row) are shown to highlight the novelty or the commonality of the target image with respect to training and synthetic sets. Memorized samples are duplicates of the training points. Spurious samples are copies of the synthetic set that do not appear in the training set. Generalized samples do not belong to either of the two sets, indicating that they are genuinely novel samples that are not generated frequently.
@@ 2 Diffusion Models and Dense Associative Memories
Given a dataset of i.i.d. samples ${\mathbf{y}}\in\mathbb{R}^{N}$ drawn from a target data distribution $p({\mathbf{y}})$, diffusion models are a class of generative models, which aims to approximate $p({\mathbf{y}})$ by placing a reversible process that maps data to noise and back. The mapping to noise (or forward process) is described by the following Itô Stochastic Differential Equation (SDE) [25]:
$\mathrm{d}{\mathbf{x}}_{t}=\mathbf{f}({\mathbf{x}}_{t},t)\mathrm{d}t+{g}(t)\mathrm{d}{\mathbf{w}}_{t}$ (1)
@@
which transforms the given data distribution (${\mathbf{x}}_{0}={\mathbf{y}}$) into a simpler one, such as an isotropic Gaussian distribution. Here, ${\mathbf{w}}_{t}$ is the standard Wiener process (or Brownian motion) and $\mathbf{f}({\mathbf{x}}_{t},t)$ denotes the drift term that guides the diffusion process, which we will assume to be zero for the most part of this paper. Meanwhile, $g(t)$ represents the diffusion coefficient that controls the noise at each time step $t\rightarrow T$. The reverse process removes the injected noise and is described as
$\mathrm{d}{\mathbf{x}}_{t}=[\mathbf{f}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})]\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (2)
@@
where $\bar{{\mathbf{w}}}_{t}$ is the standard Wiener process. To effectively solve this equation, one must reliably estimate the score $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ via parameterization of the score as a neural network $s_{\theta}({\mathbf{x}},t)$, where $\theta$ denotes neural network’s trainable parameters. This network is trained using methods for denoising score matching across multiple times steps [23, 25, 56, 57, 58].
Our theoretical arguments pertain to the idealistic setting in which this neural network is capable of learning the true empirical distribution obtained by adding noise to the training data. This limiting case is often adopted in theoretical studies of diffusion models, see [54, 59, 60, 55, 61]. From the practical perspective, this is achieved if the neural network is large (over-parametrized) and trained for a long time, an assumption that may deviate from the practical use cases. This simplification allows us to make a theoretical prediction that diffusion models must generate spurious states. It is highly non-trivial that this theoretical prediction can be extrapolated to the regime in which the size of the score-modeling neural network is finite and the amount of training data is large. The main contribution of our work is the empirical demonstration that this extrapolation is valid, and spurious states indeed exist in practically used settings.
@@
Consider the training data distribution in the variance-exploding setting of $f({\mathbf{x}}_{t},t)=0$ and $g(t)=\sigma$. In this case, the marginal probability distribution of new samples can be computed exactly as
@@
$p({\mathbf{x}}_{t},t)=\underset{{\mathbf{y}}\sim\text{data}}{\mathbb{E}}\Bigg{[}\frac{1}{(2\pi\sigma^{2}t)^{\frac{N}{2}}}\exp\Big(-\frac{\lVert{\mathbf{x}}_{t}-\mathbf{y}\rVert^{2}_{2}}{2\sigma^{2}t}\Big{missing})\Bigg{]}$ (3)
@@
Assuming the empirical distribution of the data $p(\mathbf{y})=\frac{1}{K}\sum\limits_{\mu=1}^{K}\delta^{(N)}(\mathbf{y}-\bm{\xi}^{\mu})$, where $\bm{\xi}^{\mu}$ represents an individual data point (with data size $K$), this marginal distribution can be written as

### v3

@@ Abstract
Dense Associative Memories (DenseAMs) are generalizations of Hopfield networks, which have superior information storage capacity and can store training data points (memories) at local minima of the energy landscape. When the amount of training data exceeds the critical memory storage capacity of these models, new local minima, which are different from the training data, emerge. In Associative Memory these emergent local minima are called spurious states, which hinder memory retrieval. In this work, we examine diffusion models (DMs) through the DenseAM lens, viewing their generative process as an attempt of a memory retrieval. In the small data regimes, DMs create distinct attractors for each training sample, akin to DenseAMs below the critical memory storage. As the training data size increases, they transition from memorization to generalization. We identify a critical intermediate phase, predicted by DenseAM theory – the spurious states. In generative modeling, these states are no longer negative artifacts but rather are the first signs of generative capabilities. We characterize the basins of attraction, energy landscape curvature, and computational properties of these previously overlooked states. Their existence is demonstrated across a wide range of architectures and datasets. Corresponding email: phamb@rpi.edu. The code is available from this link. denotes equal contribution.
@@ 1  Introduction
Hopfield networks are energy-based Associative Memory (AM) models, which conceptualize memories as attractor states corresponding to local minima of their energy function, and the memory retrieval as dynamical convergence towards such attractors Hopfield (1982, 1984); Amari (1972). Modern variants of such networks, called Dense Associative Memories (DenseAMs) Krotov and Hopfield (2016, 2018), have helped revitalize scientific interest in these ideas and paved the way for more sophisticated AM systems Demircigil et al. (2017); Agliari et al. (2020); Agliari and De Marzo (2020); Albanese et al. (2022); Millidge et al. (2022); Saha et al. (2023); Krotov (2023); Hoover et al. (2023a), driven primarily by their connection to the attention mechanism in Transformers Ramsauer et al. (2021); Hoover et al. (2023a) and neurobiology Krotov and Hopfield (2021); Kozachkov et al. (2025).
Simultaneously, diffusion models (DMs) Sohl-Dickstein et al. (2015) have gained popularity, due to their flexibility and accuracy in modeling various high-dimensional distributions Ho et al. (2020); Song and Ermon (2019); Song et al. (2021); Rombach et al. (2022). However, despite their effectiveness, DMs pose challenges related to privacy and security, as concerns grow about their tendency to replicate training data alongside a lack of understanding of how generalization arises in these models Somepalli et al. (2023a, b); Carlini et al. (2023); Wen et al. (2024); Jeon et al. (2024); Webster (2023). Hence, such matters emphasize the need for further understanding of memorization and generalization behaviors of DMs.
@@
Currently, much of the recent works on the memorization-to-generalization transition of DMs tackle it in a generalization-centric fashion Meehan et al. (2020); den Burg and Williams (2021); Somepalli et al. (2023a, b); Kadkhodaie et al. (2023); Ventura et al. (2024); Achilli et al. (2024); Kamb and Ganguli (2024); Ross et al. (2024); Biroli et al. (2024); Cui et al. (2025); Achilli et al. (2025). Namely, the memorization phenomenon is typically viewed as a “small side effect,” alleviated by drastically increasing the training data size Yoon et al. (2023); Gu et al. (2023). Although these studies shed some light on memorization and generalization, they do not fully explore the intermediate transition regime between these two phases. Our work adopts a complimentary approach which further ties memorization and generalization together. From the beginning, we cast the diffusion modeling pipeline into the AM framework. The training phase of diffusion modeling is conceptualized as the operation of writing the training data into the memory. The generation phase is viewed as the attempt of memory recall. This recall can be successful, resulting in the retrieval of training samples (memorization), or unsuccessful, resulting in the generation of new previously unseen samples (generalization). Consequently, this point of view allows us to apply the theory developed for AMs to the memorization-to-generalization transition in DMs.
Recent works Hoover et al. (2023b); Ambrogioni (2024); Raya and Ambrogioni (2024) have also begun establishing theoretical connections between DenseAMs and DMs, illustrating that the score function of typical DMs can be interpreted as the gradient of a DenseAM’s energy function. These prior works, however, do not discuss a cornerstone phenomenon of AMs – the emergence of spurious states – which can further bridge both DenseAMs and DMs. Historically considered as detrimental to reliable memory recall Hopfield (1982); Hopfield et al. (1983); Hopfield (1984); Amit et al. (1985); Abu-Mostafa and St. Jacques (1985), spurious states were viewed as negative artifacts of AM networks. In contrast, for generative modelling these same spurious states serve as the first signatures of generalization capabilities, thus becoming a desirable emergent phenomenon Kalaj et al. (2025).
@@
In accord with previous literature, we empirically find that DMs undergo a transition from memorization to generalization as the number of training points increases on realistic datasets. Unlike previous works, at the onset of generalization we detect the emergence of spurious patterns, which have sizable basins of attraction, but do not correspond to memorized data points. No longer negative artifacts, spurious states are now observed to be positive indicators of early generalization in DMs and they may also exist even during the full generalization stage of DMs, including Stable Diffusion Rombach et al. (2022). Lastly, we provide theoretical descriptions distinguishing these peculiar states from memorized and generalized patterns, in terms of energy landscapes. Our findings suggest a close parallel between two seemingly different systems, DMs and DenseAMs, studied by two disjoint communities. Specifically, the DenseAM’s perspective on DMs can illuminate some of the computational properties of DMs at the boundary of the memorization-generalization transition.
Figure 1: Panel (A) schematically illustrates the change in the energy landscape as the size of the training dataset is increased. In the small data regime, the model stores the training data points as local minima of the energy. When the amount of training data exceeds the model’s memory capacity, spurious patterns are formed and training data points are no longer energy minima. Subsequent increase of the training data size leads to the generalization phase, defined by the formation of continuous manifold of the low energy states. Examples of memorized, spurious, and generalized samples in their respective columns for four datasets (MNIST, FASHION-MNIST, CIFAR10, and LSUN-CHURCH) are provided in Panel (B), see Sec. (4.1) for our definitions of these three sample types. For each target image (shown on the left), its top-4 nearest neighbors from the training set (top row) and the synthetic set (bottom row) are shown to highlight the novelty and commonality of the target image with respect to its training and synthetic sets. To help highlight the novelty of spurious patterns, we provide rectangular markers guiding where the features differ in their corresponding nearest neighbors.
@@ 2  Diffusion Models and Dense Associative Memories
Motivated by Hoover et al. (2023b); Ambrogioni (2024); Raya and Ambrogioni (2024), we first provide a theoretical link between DMs and DenseAMs through the perspective of energy. Given a dataset of samples ${\mathbf{y}}\in\mathbb{R}^{N}$ drawn from a target data distribution ${\mathbf{y}}\overset{\text{i.i.d}}{\sim}p({\mathbf{y}})$, we rely on the following stochastic differential equation (SDE) which describes the forward process Song et al. (2021):
$\mathrm{d}{\mathbf{x}}_{t}=\mathbf{f}({\mathbf{x}}_{t},t)\mathrm{d}t+{g}(t)\mathrm{d}{\mathbf{w}}_{t},$ (1)
@@
which transforms $p({\mathbf{y}})$, where ${\mathbf{x}}_{0}={\mathbf{y}}$, into a simpler distribution, e.g., isotropic Gaussian. Here, ${\mathbf{w}}_{t}$ is the standard Wiener process and $\mathbf{f}({\mathbf{x}}_{t},t)$ denotes the drift term that guides the diffusion process, which we assume to be zero for most of the paper. Meanwhile, $g(t)$ denotes the diffusion coefficient controlling the noise at each time step $t\rightarrow T$. The reverse, denoising, or generative process is described as
$\mathrm{d}{\mathbf{x}}_{t}=[\mathbf{f}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})]\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t},$ (2)
@@
where $\bar{{\mathbf{w}}}_{t}$ is the standard Wiener process. To effectively solve this equation, one must reliably estimate the score $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})=-\nabla_{{\mathbf{x}}_{t}}E({\mathbf{x}}_{t},t)$ via training a neural network $s_{\theta}({\mathbf{x}}_{t},t)$, where $\theta$ denotes its set of trainable parameters, using methods for denoising score matching across multiple time steps Ho et al. (2020); Song et al. (2021); Hyvärinen (2005); Vincent (2011).
@@
Consider the training data distribution in the variance-exploding setting of $f({\mathbf{x}}_{t},t)=0$ and $g(t)=\sigma$. In this case, the marginal probability distribution of noisy samples at time $t$ can be computed exactly as
@@
$p({\mathbf{x}}_{t},t)=\underset{{\mathbf{y}}\sim\text{data}}{\mathbb{E}}\Bigg[\frac{1}{(2\pi\sigma^{2}t)^{\frac{N}{2}}}\exp\Big(-\frac{\lVert{\mathbf{x}}_{t}-\mathbf{y}\rVert^{2}_{2}}{2\sigma^{2}t}\Big)\Bigg].$ (3)
@@
Assuming the empirical distribution of the data $p(\mathbf{y})=\frac{1}{K}\sum\limits_{\mu=1}^{K}\delta^{(N)}(\mathbf{y}-\bm{\xi}^{\mu})$, where $\bm{\xi}^{\mu}$ represents an individual data point (belonging to a dataset with training data size $K$), this marginal distribution can be written as
$p({\mathbf{x}}_{t},t)=\frac{1}{K}\sum\limits_{\mu=1}^{K}\frac{1}{(2\pi\sigma^{2}t)^{\frac{N}{2}}}\exp\Big({-\frac{\lVert{\mathbf{x}}_{t}-\bm{\xi}^{\mu}\rVert^{2}_{2}}{2\sigma^{2}t}}\Big)\overset{\text{def}}{\equiv}\exp\Bigg(-{\frac{E^{\text{DM}}({\mathbf{x}}_{t},t)}{2\sigma^{2}t}}\Bigg),$ (4)
where we also defined the energy $E^{\mathrm{DM}}$ of the DM, up to $\mathbf{x}$}–independent terms, to be
$E^{\text{DM}}({\mathbf{x}}_{t},t)=-2\sigma^{2}t\log\bigg[\sum\limits_{\mu=1}^{K}\exp\Big(-\frac{\lVert{\mathbf{x}}_{t}-\bm{\xi}^{\mu}\rVert^{2}_{2}}{2\sigma^{2}t}\Big)\bigg].$ (5)
As already observed in Ambrogioni (2024), Eq. (5) is closely related to a commonly studied DenseAM’s model, see for example Saha et al. (2023):
$E^{\text{AM}}({\mathbf{x}})=-\beta^{-1}\log\bigg[\sum\limits_{\mu=1}^{K}\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{\mu}\rVert^{2}_{2}\Big)\bigg].$ (6)
A direct comparison of Eqs. (5) and (6) highlights a close connection between these two seemingly different frameworks. Specifically, the data points $\bm{\xi}^{\mu}$ in the DM framework play the role of memories in the AM formulation. The variance of the noise added during the forward process (1) plays the role of the effective temperature $\beta^{-1}$ (which controls the separation of memories), while the reverse process (2) corresponds to memory retrieval dynamics Hoover et al. (2023b); Ambrogioni (2024).
For a single data point $\bm{\xi}^{1}$ or $K=1$, Eq. (6) has a single local minimum at that stored pattern, where the shape of this energy landscape is independent of the value of $\beta$. This is an example of a memorized state. In contrast, for $K=2$, there are several possibilities. The energy landscape has two local minima when $\beta\rightarrow\infty$ is large. But, for finite values of $\beta$, there exist configurations of patterns such that Eq. (6) has only one local minimum:
$\bm{\eta}=\operatorname*{arg\,min}_{{\mathbf{x}}}E^{\text{AM}}({\mathbf{x}}),$ (7)
such that $\bm{\eta}\neq\bm{\xi}^{1}$ and $\bm{\eta}\neq\bm{\xi}^{2}$. This is an example of a spurious state. Such states are ubiquitous in AM models, when the amount of training data (memories) crosses the critical memory storage capacity of the model (which depends on the correlations between the training data points). Thus, the following question can be posed: Is it possible to observe spurious states in DMs trained using score matching objective when the energy landscape is modeled by the score-predicting neural network?
3  Discussion
We argue below that the answer to the above question is positive, and the spurious states indeed emerge, as DMs progress from memorization to generalization as the training data size $K$ increases. As we observed numerically in Fig. (4A), when $K$ is small, DMs initially memorize their training data points and store them as local minima of their energy function. In the large data regime, a different phase appears where a sufficient increase of $K$ fosters the creation of new attractor states corresponding to the manifold of the generated samples. Spurious states appear at the boundary of this transition and correspond to emergent attractor states of the energy function, absent in the training set, but at the same time, having distinct basins of attraction around them. Subsequent increases of $K$ leads to a decrease of the number of memorized and spurious states, and the creation of generalized states, points that occupy the continuous manifold. This sequence of transitions is illustrated in Fig. (1A).
In order to empirically establish this sequence of transitions, we have trained a set of conventional diffusion models on datasets of gradually increasing sizes. Each of the trained models is characterized by its training set $\mathsf{S}$, and a synthetic set of generated samples drawn from this model $\mathsf{S}^{\prime}$ (the synthetic set is assumed to be very large). A sample $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}$ is memorized if it has duplicates in both training and synthetic sets. This is clear from the three universality classes shown in Fig. (1A). Since memorized samples have a basin of attraction around them, the same sample should appear in the synthetic set more than once, and that generated sample should be a copy of the training data point. A sample $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}$ is spurious if it has duplicates in the synthetic set, but not in the training set. Again, this is clear from Fig. (1A): spurious samples have a well-defined basin of attraction around them, but the bottom of that basin corresponds to a newly generated sample – not the training example. Finally, the generalized sample $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}$ has no duplicates in either the training or synthetic sets. Generalized samples occupy flat manifolds in the configuration space. For this reason, every generation produces a slightly different sample resulting in absence of duplicates in $\mathsf{S}^{\prime}$. Additional details pertaining to these detection metrics are given in section 4.
Examples of generated samples belonging to these three categories are shown in Fig. (1B). For each sample, the top-4 nearest neighbors from the training set (top row) and synthetic set (bottom row) are shown. It is instructive to visually examine some of the samples classified as spurious. For instance, the first spurious sample in Fig. (1B) (the model was trained on CIFAR-10 dataset Krizhevsky et al. (2014)) is a plane that has a large wing (highlighted by the cyan frame), which is absent in the training set, but has an exact duplicate in the synthetic set. Another interesting example is shown in the 4-th row (the model was trained on FASHION-MNIST dataset Xiao et al. (2017)). There are many examples of shirts in the training set with either two sleeves, or no sleeves. The generated sample shows only one sleeve. This sample has exact duplicates in the synthetic set, indicating the existence of a well-defined basin of attraction around it, but does not appear in the training set. These examples illustrate the general phenomenon, which is predicted by AM perspective and is clearly observed empirically for conventional diffusion models trained with the score matching objective.
The abundance of memorized, spurious, and generalized samples is quantified as a function of training data set size in Fig. (4A). As training set gets larger, the amount of memorized samples decreases, while the amount of generalized samples increases. The spurious states appear as a peak at the boundary of this transition.
The distinct nature of these three different states is further substantiated by the hierarchy of the log-volume of their basins of attraction in Fig. (4B) and their energy curvature spectra in Fig. (5A). Regarding the basin volume, memorized samples consistently exhibit the largest log-volume, followed by spurious samples, while generalized samples have near-zero volume. This is in accord with the intuitive expectation for the energy profiles in Fig. (1A). Memorized and spurious samples should have substantial basins of attraction around them, while generalized samples should have almost vanishing basins, since they occupy an almost flat energy manifold. The energy curvature spectral analysis reveals a similar finding: as the training data size $K$ increases: the memorized and spurious samples tend to have higher curvature (indicated by their fewer near-zero singular values) compared to the generalized samples.
Initially, our studies focused on relatively small scale datasets: MNIST Deng (2012), FASHION-MNIST Xiao et al. (2017), CIFAR10 Krizhevsky et al. (2014), and LSUN-CHURCH Yu et al. (2015). Focusing on these settings allowed us to train a dense grid of models with varying size of the training set and carefully examine the existence of duplicates in the training and synthetic sets. Such a careful analysis would be intractable for large-scale real world scenarios such as Stable Diffusion Rombach et al. (2022): training these models requires substantial compute resources, the training set is large, and the synthetic set needs to be even larger in order to reliably perform the analysis similar to what was described above. At the same time, the analysis based on the curvature of the energy landscape around generated samples is more feasible.
Thus, we have identified several candidate samples from Stable Diffusion that fall into three categories: sharp energy curvature, intermediate energy curvature, and almost flat energy. Examples of these samples together with their singular values of the score function Jacobian are shown in Fig. (5B). For each sample a language prompt was given to the model and the generated samples are shown together with the training image corresponding to that prompt. High curvature samples (candidate memorized states) are exact duplicates of the training image. Intermediate curvature samples (candidate spurious states) are somewhat similar to the training image, but deviate from it in certain aspects (color and pattern of the rug, color of the curtains). Finally, small curvature samples (candidate generalized states) demonstrate a greater degree of variability and larger deviations from the training image. Qualitatively, these results align with our primary findings on the carefully studied small dataset models.